What is eqivalent charge when we add same resistance in series.
1
Expert's answer
2020-12-25T14:07:44-0500

We will solve this problem by this way:

Let`s find eqivalent charge:

q1="I\times t=\frac{U}{R1}\times t" , q1-charge, I-current, t- spent time to flow current, R1="n\times R" -total resistance, U-voltage.

If we add a same resistor to circuit, then:

(total resistance) R2="n\times R" +"R", this means that total resistance increases as we add new resistors and also it means :

q2="I\times t=\frac{U}{n*R+R}\times t". In this case, charge decreases as resistance increases

q2"\lt"q1 because R1"\lt" R2